OpenStudy (anonymous):

Write an equation in point-slop form of the line through joint J(4, 1) with slope -4. please help ?? :))

OpenStudy (compassionate):

y - y1 = m(x - x1) Where my is the slope and y1 and x1 are your points.

OpenStudy (anonymous):

ooo, okay i get it now. haha, thank you so much !!

OpenStudy (compassionate):

No problem. Remember The forms of: STANDARD Ax + Bx = C POINT-SLOPE y - y1 = m(x - x1) SLOPE INTERCEPT FORM Y = mx + b If you are EVER given something like a slope and points, and asked to graph it, first put it in POINT SLOPE FORM then solve for y, that will give you y = mx + b, and you can graph it then.